OpenStudy (anonymous):

200 is the ? th term of 24, 35, 46, 57.......

OpenStudy (cwrw238):

this is an arithmetic series with first term a1 = 24 and common difference = 11 and nth term = a1 + (n - 1)*d so plug in n = 200 and values of a1 and d and you've got it

OpenStudy (cwrw238):

* correction nth term = 200 then solve for n: so we have 200 = 24 + (n - 1)* 11 solve this for n

OpenStudy (anonymous):

I don't think I asked my question correctly. I have the answer that 200 is the 17th term of 24, 35, 46, 57...... I just don't know how to work it out! HELP PLEASE!